### v1

@@ Appendix B Future Prospects: New Periodic Variables among the non-HNEV objects
Optical spectroscopic observation of PN G030.8+03.4 (one of the objects showing outbursts in the ZTF light curve, see Figure 3) was performed by DBSP on the same date with the same instrument configuration as WeSb 1 (see Section 4.2.1). The same reduction procedure was also followed. Figure C1 shows the spectrum for the object. The unusual feature in the spectra are the strong Ca II (H, K, and triplet) emission lines, which are not usually seen in PNe. The spectrum also lacks the [O III] emission lines. Considering the light curve and lack of nebula, the spectrum is indicative of either a symbiotic system or a CV. We also see prominent [S II] 6716, 6731 Å emissions in the spectrum. In CVs, such forbidden lines appear when the system has sufficiently evolved since the nova eruption (old novae, see for example the discussion in Inight et al. 2023). Further observations with better signal-to-noise are needed to confirm the nature of the object, but it is unlikely to be a PN.
Appendix D Misidentified nucleus of WeSb 1?
@@
We now briefly discuss the possibility that this stellar source is not the true nucleus of WeSb 1. With its high-density compact emission and unusual light curve, the object is very unique. The odds of a chance appearance of such a rare system at the center of a 3′ nebula is extremely small. Nevertheless, we performed a brief search for other candidate WeSb 1 central stars in the vicinity of the current nucleus. We queried Gaia catalog within 45′′ of the present nucleus (about half the radius of the nebula). The Gaia parallax distances to most of these 26 sources are much larger than the approximate nebular distance (say for example as estimated in Frew et al. 2016), albeit determined with very low parallax-over-errors (POEs). We then manually set the distances of the stars (with low POE) to the nebular distance and compared their position in the Gaia CMD. At this distance, the current nucleus is the brightest, though not the bluest. However, none of the other stars showed immediate promise in becoming a candidate for WeSb 1 nucleus. Furthermore, within this radius, only three objects (other than the current nucleus) have a GALEX NUV counterpart. Two objects were rejected for being a nearby star and for geometrical reasons, respectively. The remaining star (closest to the current nucleus in the southeast direction) seemed the most promising alternative. However, the Gaia estimated distance is 51251226 pc, which is just marginally consistent with the nebular distance. Manually placing it at the nebular distance, however, makes it much fainter and broadly consistent with a main-sequence star in the Gaia CMD. Overall, with our preliminary check, we do not find any obvious evidence to discard the current source being the nucleus of WeSb 1, though further confirmatory studies are needed. Even in such a remote possibility of misidentification, the observed system properties are significantly interesting and unique in themselves, even without the nebular association.
Appendix E HET Spectrum of WeSb 1
@@
The one-dimensional HET spectrum of the central star of WeSb 1 is shown in the top panel of Figure E1. The same features as in DBSP are recovered, in line with expectations. In the bottom two panels of the same figure, we present the comparison of the HET spectrum with the synthetic Coelho and the reference F5 IV star HD 155078 spectra (same as Figure 9, with the comparison spectra appropriately down-resolved to match the resolution of HET observation). As with DBSP, a good match is observed.
Appendix F Possible evolutionary stage of the WeSb 1 CSPN
@@
As discussed in section 4.4, the assumption that the F-type companion is truly in the subgiant phase gives a mass constraint of . With the PN still visible (thus having evolved off the main sequence just recently), this readily gives a similar mass constraint on the PN progenitor. The current properties of the CSPN could not be well-constrained due to the lack of data points at smaller wavelengths. We take the parameters of the example ‘fit’ in the bottom panel Figure 10 and examine its position in the stellar evolutionary tracks. The comparison with MIST models is presented in Figure F1 and we indeed see a reasonable agreement.
@@
Appendix G Known Periodic Variables
We present the ZTF light curves (both raw and phase-folded) and the periodograms for the known periodic CSPNe recovered in the HNEV sample in Figure G1.
Appendix H List of the short-timescale variables
We provide the list of the short-timescale HNEV variables in Table H. The non-obvious column names are – PNStat: the PN status as on the current HASH catalog; PNRad: the PN radius as in the CW21 which are derived from those in the HASH catalog; Rel: the reliability score (in the range of 0 – 1) assigned to the central star identification in CW21.

### v2

@@ Appendix D Future Prospects: New Periodic Variables among the non-HNEV objects
Optical spectroscopic observation of PN G030.8+03.4 (one of the objects showing outbursts in the ZTF light curve, see Figure 3) was performed by DBSP on the same date with the same instrument configuration as WeSb 1 (see Section 4.2.1). The same reduction procedure was also followed. Figure E1 shows the spectrum for the object. The unusual feature in the spectra are the strong Ca II (H, K, and triplet) emission lines, which are not usually seen in PNe. The spectrum also lacks the [O III] emission lines. Considering the light curve and lack of nebula, the spectrum is indicative of either a symbiotic system or a CV. We also see prominent [S II] 6716, 6731 Å emissions in the spectrum. In CVs, such forbidden lines appear when the system has sufficiently evolved since the nova eruption (old novae, see for example the discussion in Inight et al. 2023). Further observations with better signal-to-noise are needed to confirm the nature of the object, but it is unlikely to be a PN.
Appendix F Misidentified nucleus of WeSb 1 or Chance Alignment?
@@
We now briefly discuss the possibility that this stellar source is not the true nucleus of WeSb 1. With its high-density compact emission and unusual light curve, the object is very unique. The odds of a chance appearance of such a rare system at the center of a 3′ nebula is extremely small. Nevertheless, we performed a brief search for other candidate WeSb 1 central stars in the vicinity of the current nucleus. We queried Gaia catalog within 45′′ of the present nucleus (about half the radius of the nebula). The Gaia parallax distances to most of these 26 sources are much larger than the approximate nebular distance (say for example as estimated in Frew et al. 2016), albeit determined with very low parallax-over-errors (POEs). We then manually set the distances of the stars (with low POE) to the nebular distance and compared their position in the Gaia CMD. At this distance, the current nucleus is the brightest, though not the bluest. However, none of the other stars showed immediate promise in becoming a candidate for WeSb 1 nucleus. Furthermore, within this radius, only three objects (other than the current nucleus) have a GALEX NUV counterpart. Two objects were rejected for being a nearby star and for geometrical reasons, respectively. The remaining star (closest to the current nucleus in the southeast direction) seemed the most promising alternative. However, the Gaia estimated distance is 51251226 pc, which is just marginally consistent with the nebular distance. Manually placing it at the nebular distance, however, makes it intrinsically much fainter and broadly consistent with a main-sequence star in the Gaia CMD. Overall, with our preliminary check, we do not find any obvious evidence to discard the current source being the nucleus of WeSb 1, though further confirmatory studies are needed. Even in such a remote possibility of misidentification, the observed system properties are significantly interesting and unique in themselves, even without the nebular association.
We now consider the possibility where the candidate companion star is a chance alignment with the GALEX source, which is the true CSPN. In this scenario, the age estimation of the PN will be erroneous. We compute the probability of such a chance alignment. The distance between the GALEX and Gaia coordinates is 0.54 arcseconds. The G magnitude of the candidate central star is 14.77. We take a radius of 7 arcminutes and find 65 sources as bright as the current star (). Assuming an uniform space density, a chance alignment of such a star with the GALEX coordinate is , which is quite low. Additionally, the Gaia parallax distance, based on the optical source, is compatible to the nebular distance estimated from H surface brightness and radius relation (see Section 4). This strongly suggests a true association between the optical and the UV sources and the nebula.
A second possibility is the chance alignment of the GALEX source with the optical source, the latter being the true CSPN. This is even less likely as the GALEX source density is sparser (25 sources within 7 arcminutes). Also, a hot object (thus a GALEX source) is expected as the star of the candidate companion’s property cannot produce high excitation emission lines like [O III].
Appendix G HET Spectrum of WeSb 1
@@
The one-dimensional HET spectrum of the nucleus of WeSb 1 is shown in the top panel of Figure G1. The same features as in DBSP are recovered, in line with expectations. In the bottom two panels of the same figure, we present the comparison of the HET spectrum with the synthetic Coelho and the reference F5 IV star HD 155078 spectra (same as Figure 9, with the comparison spectra appropriately down-resolved to match the resolution of HET observation). As with DBSP, a good match is observed.
Appendix H Possible evolutionary stage of the WeSb 1 CSPN
@@
As discussed in section 4.4, the assumption that the F-type companion is truly in the subgiant phase gives a mass constraint of . With the PN still visible (thus having evolved off the main sequence just recently), this readily gives a similar mass constraint on the PN progenitor. The current properties of the CSPN could not be well-constrained due to the lack of data points at smaller wavelengths. We take the parameters of the example ‘fit’ in the bottom panel Figure 10 and examine its position in the stellar evolutionary tracks. The comparison with MIST models is presented in Figure H1 and we indeed see a reasonable agreement.
@@
Appendix I Known Periodic Variables
We present the ZTF light curves (both raw and phase-folded) and the periodograms for the known periodic CSPNe recovered in the HNEV sample in Figure I1.
Appendix J List of the short-timescale variables
We provide the list of the short-timescale HNEV variables in Table J. The non-obvious column names are – PNStat: the PN status as on the current HASH catalog; PNRad: the PN radius as in the CW21 which are derived from those in the HASH catalog; Rel: the reliability score (in the range of 0 – 1) assigned to the central star identification in CW21.